# thanks: deceasedthanks: deceased

### v1

@@ VI Summary
We conducted measurements of the $K^{-}+{{}^{3}\rm He}$ reactions resulting in mesonic final states, namely $\pi^{\mp}\Sigma^{\pm}p+n^{\prime}$, $\pi^{+}\Lambda n+n^{\prime}$, and $\pi^{-}\Lambda p+p^{\prime}$, using an incident $K^{-}$ momentum of $1~{}{\rm GeV}/c$. To investigate the mesonic decay of the $\bar{K}\!N\!N$ state, we measured the two-dimensional distributions of the invariant mass and momentum transfer of the $\pi Y\!N$ systems. The observed two-dimensional distributions of the $\pi Y\!N$ systems were well described by the model functions, which included contributions from $\bar{K}\!N\!N$ production and quasi-free $\bar{K}$ absorption processes as a result of reactions given in Eqs.(1) and (2). These model functions are essentially the same as those introduced in Ref. [2] to explain the $K^{-}+{{}^{3}\rm He}\to\Lambda p+n^{\prime}$ reaction by incorporating appropriate phase-space distributions.
@@
The fit suggests that the mesonic decay is the dominant decay branch of the $\bar{K}\!N\!N$. It was also suggested that the branching ratio of $\bar{K}\!N\!N\to\pi^{+}\Lambda n$ is similar to that of $\bar{K}\!N\!N\to\pi^{\mp}\Sigma^{\pm}p$, indicating that the mesonic $\bar{K}N$ absorption process in the $I_{\bar{K}N}=1$ channel plays an important role similar to the $I_{\bar{K}\!N}=0$ channel. The significant fraction of decay into the $I_{\bar{K}N}=1$ channel, specifically the $\pi\Lambda N$ decay, would be a major contributing factor to the broader decay width of the $\bar{K}\!N\!N$ resonance compared to that of the $\Lambda(1405)$. The presence of the $I_{\bar{K}\!N}=1$ cusp in the $\pi Y$ subsystem (Fig. 13) suggests that the real part of the $I_{\bar{K}\!N}=1$ interaction is also attractive, although it is not sufficiently strong to form a bound state in this channel.
@@
The branching ratio of the $\bar{K}\!N\!N$ between the non-mesonic and mesonic channels provides valuable information regarding the size of the system. This is because the strength of mesonic decay is directly proportional to the nucleon density. In contrast, the strength of non-mesonic decay is proportional to the square of the nucleon density, as it requires simultaneous interaction of the antikaon with two nucleons. Based on the present study, the $Br(YN)/Br(\pi Y\!N)$ ratio is estimated to be approximately 1/10 obtained as integration over all mass regions. However, we should note that the ratio becomes smaller if the integration range is up to the $\bar{K}$ binding threshold. A more precise analysis, especially using a more realistic model function, is necessary to compare the experimental results with the theoretical calculations. This analysis would provide more information about the size of the system.
@@
Further study of the multi-nucleon absorption of the $K^{-}$ beam could provide valuable insights into the clustering mechanism within the nucleus. The data show that the kaon’s direct two-nucleon absorption (2NA) process is predominantly observed in the ”$pn$” pair. Conversely, almost no reaction is observed in the ”$pp$” pair. This difference can intuitively indicate a stronger coupling or clustering of the ($pn$) pair compared to the ($pp$) pair in ${{}^{3}\rm He}$.
@@
It is crucial to conduct new experiments with enhanced acceptance coverage and improved neutron detection efficiency to understand these issues better and investigate the production mechanism of kaonic nuclei. These advancements will enable us to gather more comprehensive data and provide valuable insights into the phenomena under investigation. By increasing the experimental capabilities, we can further explore the properties and behavior of kaonic nuclei, shed light on their production mechanisms, and contribute to a more thorough understanding of these fascinating systems [26].
@@ Acknowledgements.
The authors are grateful to the staff members of J-PARC/KEK for their extensive efforts, especially in the stable operation of the facility. We are also grateful to the contributions of Professors D. Jido, T. Sekihara, Y. Akaishi, A. Dote, T. Harada, O. Morimatsu, J. Yamagata, and Dr. K. Suzuki. This work is partly supported by MEXT Grants-in-Aid 14102005, 17070007, 24105003, 26287057, 26800158, 17K05481, 18H05402, 20K04006, 21K13952, and 22H04917. Part of this work is supported by the EU STRONG-2020 project (Grant Agreement No. 824093) and by the EXOTICA project of the Minstero degli Affari Esteri e della Cooperazione Internazionale, PO22MO03.

### v2

@@ VI Summary
We conducted measurements of the $K^{-}+{{}^{3}\rm He}$ reactions resulting in mesonic final states—namely, $\pi^{\mp}\Sigma^{\pm}p+n^{\prime}$, $\pi^{+}\Lambda n+n^{\prime}$, and $\pi^{-}\Lambda p+p^{\prime}$—using an incident $K^{-}$ momentum of $1~{}{\rm GeV}/c$. To investigate the mesonic decay of the $\bar{K}\!N\!N$ state, we measured the two-dimensional distributions of the invariant mass and momentum transfer of the $\pi Y\!N$ systems. The observed two-dimensional distributions of the $\pi Y\!N$ systems were well described by the model functions, which included contributions from $\bar{K}\!N\!N$ production and quasi-free $\bar{K}$ absorption processes as a result of the reactions given in Eqs. (1) and (2). These model functions are essentially the same as those introduced in Ref. [2] to explain the $K^{-}+{{}^{3}\rm He}\to\Lambda p+n^{\prime}$ reaction by incorporating appropriate phase-space distributions.
@@
The fitting results suggest that the mesonic decay is the dominant decay branch of the $\bar{K}\!N\!N$. The results also suggest that the branching ratio of $\bar{K}\!N\!N\to\pi^{+}\Lambda n$ is similar to that of $\bar{K}\!N\!N\to\pi^{\mp}\Sigma^{\pm}p$, indicating that the mesonic $\bar{K}N$ absorption process in the $I_{\bar{K}N}=1$ channel plays an important role, similar to the process in the $I_{\bar{K}\!N}=0$ channel.The significant fraction of decay into the $I_{\bar{K}N}=1$ channel—specifically, the $\pi\Lambda N$ decay—would be a major factor contributing to the broader decay width of the $\bar{K}\!N\!N$ resonance compared with that of the $\Lambda(1405)$. The presence of the $I_{\bar{K}\!N}=1$ cusp in the $\pi Y$ subsystem (Fig. 13) suggests that the real part of the $I_{\bar{K}\!N}=1$ interaction is also attractive, although it is not sufficiently strong to form a bound state in this channel.
@@
The branching ratio for the $\bar{K}\!N\!N$ between the non-mesonic and mesonic channels provides valuable information regarding the size of the system because the strength of mesonic decay is directly proportional to the nucleon density. By contrast, the strength of non-mesonic decay is proportional to the square of the nucleon density because it requires simultaneous interaction of the antikaon with two nucleons. On the basis of the present study, the $Br(YN)/Br(\pi Y\!N)$ ratio is estimated to be approximately 1/10, as obtained via integration over all mass regions. However, we note that the ratio becomes smaller if the integration range is limited to the $\bar{K}$ binding threshold. A more precise analysis, especially one using a more realistic model function, is necessary to compare the experimental results with the theoretical calculations. This analysis would provide more information about the size of the system.
@@
Further study of the multi-nucleon absorption of the $K^{-}$ beam could provide valuable insights into the clustering mechanism within the nucleus. The data show that the kaon’s direct two-nucleon absorption process is predominantly observed in the ”$pn$” pair. Conversely, almost no reaction is observed in the ”$pp$” pair. This difference can intuitively indicate a stronger coupling or clustering of the ($pn$) pair compared with that of the ($pp$) pair in ${{}^{3}\rm He}$.
@@
It is crucial to conduct new experiments with enhanced acceptance coverage and improved neutron detection efficiency to understand these issues better and investigate the production mechanism for kaonic nuclei. These advancements will enable us to gather more comprehensive data and will provide valuable insights into the phenomena under investigation. By increasing the experimental capabilities, we can further explore the properties and behavior of kaonic nuclei, shed light on their production mechanisms, and contribute to a more thorough understanding of these fascinating systems [26].
@@ Acknowledgements.
The authors are grateful to the staff members of J-PARC/KEK for their extensive efforts, especially in the stable operation of the facility. We are also grateful for the contributions of Professors D. Jido, T. Sekihara, Y. Akaishi, A. Dote, T. Harada, O. Morimatsu, and J. Yamagata and Dr. K. Suzuki. This work is partly supported by MEXT Grants-in-Aid 14102005, 17070007, 24105003, 26287057, 26800158, 17K05481, 18H05402, 20K04006, 21K13952, 22H04917, 22H04940, and 24H00029. Part of this work is supported by the EU STRONG-2020 project (Grant Agreement No. 824093) and by the EXOTICA project of the Minstero degli Affari Esteri e della Cooperazione Internazionale, PO22MO03.